# Fine-Tuning Discrete Diffusion Models via Reward Optimization with Applications to DNA and Protein Design

### v1

@@ 5.1 Proof Sketch of Theorem 1
$\mathbb{E}_{x_{t:T}\sim P^{\theta^{\star}}}\left[r(x_{T})-\alpha\int_{s=t}^{T}\sum_{y\neq x_{s}}\left\{Q^{\theta^{\star}}_{x_{s},y}(s)-Q^{\theta_{\mathrm{pre}}}_{x_{s},y}(s)+Q^{\theta^{\star}}_{x_{s},y}(s)\log\frac{Q^{\theta^{\star}}_{x_{s},y}(s)}{Q^{\theta_{\mathrm{pre}}}_{x_{s},y}(s)}\right\}ds\mid x_{t}=x\right].$
This represents the expected return when starting from state $x$ at time $t$ and following the optimal policy. Once the optimal value function is defined, the optimal generator can be expressed in terms of this value function, as shown below. This is derived from the Hamilton-Jacobi-Bellman (HJB) equation in CTMC.
@@
For $x\neq y$ ($x,y\in\mathcal{X}$), we have
$Q^{\theta^{\star}}_{x,y}(t)=Q^{\theta_{\mathrm{pre}}}_{x,y}(t)\exp(\{V_{t}(y)-V_{t}(x)\}/\alpha).$
@@
$\exp(V_{t}(x)/\alpha)=\mathbb{E}_{x_{t:T}\sim P^{\theta_{\mathrm{pre}}}}[\exp(r(x_{T})/\alpha)|x_{t}=x]$
@@
The marginal distribution at time $t$ by (4), $p^{\star}_{t}\in\Delta(\mathcal{X})$, is proportional to
$\exp(V_{t}(\cdot)/\alpha)p^{\mathrm{pre}}_{t}(\cdot).$
where $p^{\mathrm{pre}}_{t}\in\Delta(\mathcal{X})$ is a marginal distribution induced by pretrained model at $t$.
@@ 5.2 Relation to Classifier Guidance for Discrete Diffusion Models
$\exp(V_{t}(x)/\alpha)=\mathop{\mathrm{argmin}}_{g:\mathcal{X}\to\mathbb{R}}\mathbb{E}_{x_{T}\sim P^{\theta_{\mathrm{pre}}}}[\{\exp(r(x_{T})/\alpha)-g(x_{t})\}^{2}].$
In practice, while we can’t calculate the exact expectation, we can still replace it with its empirical analog. Alternatively, we can approximate it by using a map from $x_{t}$ to $x_{0}$ in pretrained models following DPS (Chung et al., 2022) or reconstruction guidance (Ho et al., 2022).
@@
$Q^{\theta^{\star}}_{x,y}(t)=Q^{\theta_{\mathrm{pre}}}_{x,y}(t)\times p_{t}(z|y)/p_{t}(z|x),\quad p_{t}(z|x_{t}):=\mathbb{E}_{x_{t:T}\sim P^{\theta_{\mathrm{pre}}}}[p(z|x_{T})\mid x_{t}].$
Thus, we can rederive the formula in Nisonoff et al. (2024). Here, we also note that this type of result is commonly referred to as the Doob transform in the literature on stochastic processes (Levin and Peres, 2017, Section 17).
@@ 6 Experiments
Our experiments focus on the design of regulatory DNA sequences for enhancer activity and protein sequences for stability. Our results include comprehensive evaluations, highlighting the ability of DRAKES to produce natural-like sequences while effectively optimizing the desired properties.
@@ 6.1 Baselines
We compare DRAKES against several baseline methods discussed in Section 2, which we summarize below with further details in Appendix D.1.
@@
  • Classifier-free Guidance (CFG) (Ho and Salimans, 2022). CFG is trained on labeled datasets with the measured attributes we aim to optimize.
@@ 6.2 Regulatory DNA Sequence Design
Dataset and settings. We experiment on a publicly available large-scale enhancer dataset (Gosai et al., 2023), which measures the enhancer activity of $\sim$700k DNA sequences (200-bp length) in human cell lines using massively parallel reporter assays (MPRAs), where the expression driven by each sequence is measured. We pretrain the masked discrete diffusion model (Sahoo et al., 2024) on all the sequences. We then split the dataset and train two reward oracles (one for finetuning and one for evaluation) on each subset, using the Enformer (Avsec et al., 2021) architecture to predict the activity level in the HepG2 cell line. These datasets and reward models are widely used in the literature on computational enhancer design (Lal et al., 2024; Uehara et al., 2024; Sarkar et al., 2024). Detailed information about the pretrained model and reward oracles is in Appendix D.2.
@@
  • Predicted activity based on the evaluation reward oracle (Pred-Activity). We predict the enhancer activity level in the HepG2 cell line using the reward oracle trained on the evaluation subset. Note that the diffusion models are fine-tuned (or guided) with the oracle trained on a different subset of the data, splitting based on chromosome following conventions (but in the same cell lines) (Lal et al., 2024).
@@
  • Log-likelihood of sequences (Log-Lik). We calculate the log-likelihood of the generated sequences with respect to the pretrained model to measure how likely the sequences are to be natural-like. Models that over-optimize the reward oracle generate out-of-distribution sequences and would have a low likelihood to the pretrained model. The likelihood is calculated using the ELBO of the discrete diffusion model in  Sahoo et al. (2024).
Results. DRAKES generates sequences with high predicted activity in the HepG2 cell line, as robustly measured by Pred-Activity and ATAC-Acc (Table 1). The generated sequences closely resemble natural enhancers, as indicated by high 3-mer and JASPAR motif correlations, and a similar likelihood to the pretrained model. These highlight DRAKES’s effectiveness in generating plausible high-activity enhancer sequences. Notably, while DRAKES, without KL regularization achieves higher Pred-Activity, this can be attributed to over-optimization. Despite splitting the data for fine-tuning and evaluation, the sequences remain highly similar due to many analogous regions within each chromosome. However, when evaluated with an independent activity oracle, ATAC-Acc, DRAKES  demonstrates superior performance while maintaining higher correlations and log likelihood.
@@
Method Pred-Activity (median) $\uparrow$ ATAC-Acc $\uparrow$ (%) 3-mer Corr $\uparrow$ JASPAR Corr $\uparrow$ Log-Lik (median) $\uparrow$ Pretrained 0.17(0.04) 1.5(0.2) -0.061(0.034) 0.249(0.015) -261(0.6) CG 3.30(0.00) 0.0(0.0) -0.065(0.001) 0.212(0.035) -266(0.6) SMC 4.15(0.33) 39.9(8.7) 0.840(0.045) 0.756(0.068) -259(2.5) TDS 4.64(0.21) 45.3(16.4) 0.848(0.008) 0.846(0.044) -257(1.5) CFG 5.04(0.06) 92.1(0.9) 0.746(0.001) 0.864(0.011) -265(0.6) DRAKES w/o KL 6.44(0.04) 82.5(2.8) 0.307(0.001) 0.557(0.015) -281(0.6) DRAKES 5.61(0.07) 92.5(0.6) 0.887(0.002) 0.911(0.002) -264(0.6)
@@ 6.3 Protein Sequence Design: Optimizing Stability in Inverse Folding Model
Dataset and settings. First, we pretrained an inverse folding model based on the diffusion model (Campbell et al., 2024) and the ProteinMPNN (Dauparas et al., 2022) architecture, using the PDB training set from Dauparas et al. (2022). Next, we trained the reward oracles using a different large-scale protein stability dataset, Megascale (Tsuboyama et al., 2023), which includes stability measurements (i.e., the Gibbs free energy change) for $\sim$1.8M sequence variants from 983 natural and designed domains. Following dataset curation and a train-validation-test splitting procedure from Widatalla et al. (2024) (which leads to $\sim$0.5M sequences on 333 domains) and using the ProteinMPNN architecture, we constructed two reward oracles – one for fine-tuning and one for evaluation, that predict stability from the protein sequence and wild-type conformation. Detailed information on the pretrained model and reward oracles is in Appendix D.3.
@@
Results. For inverse protein folding, DRAKES generates high-stability protein sequences capable of folding into the conditioned structure (Table 2). It achieves the highest Pred-ddG among all methods, while maintaining a similar success rate of inverse folding (measured by %(scRMSD$<2$), the percentage of scRMSD smaller than 2) as the pretrained model. Considering both factors, DRAKES significantly outperforms all baseline methods in terms of overall success rate. Note that CFG does not work well for protein sequence design due to limited labeled data, as Megascale includes only a few hundred backbones, making generalization difficult. This is expected, as we mention in Section 2. Further details are provided in Appendix D.1.
@@
For validation purposes, we calculate the stability (i.e., Gibbs free energy) of the generated sequences using physics-based simulations (PyRosetta (Chaudhury et al., 2010)) for wild-type protein backbone structures in Figure 2, following (Widatalla et al., 2024). Although all models are conditioned on the same set of protein backbones, different sets of sequences generated by generative methods can lead to significant differences in side chain interactions, which affect folding energies. The results demonstrate that sequences generated by our algorithms are more stable in this in silico validation compared to other baseline methods. For additional results, refer to Figure 6 in Appendix D.3.
@@ 7 Conclusions
Appendix A Potential Limitations
We have formulated the RL problem, (3), in the context of CTMC. The proposed algorithm in our paper to solve this problem requires reward models to be differentiable. Since differentiable models are necessary when working with experimental offline data, this assumption is not overly restrictive. Moreover, many state-of-the-art sequence models mapping sequences to functions in biology are available today, such as enformer borozi. In cases where creating differentiable models is challenging, we recommend using PPO-based algorithms (Black et al., 2023; Fan et al., 2023) or reward-weighted MLE (Uehara et al., 2024, Section 6.1).
Appendix B Proof of Theorems
B.1 Preparation
We prepare two well-known theorems in CTMC for the pedagogic purpose. For example, refer to Yin and Zhang (2012) for the more detailed proof. In these theorems, we suppose we have the CTMC:
@@ B.1 Preparation
Here, we prove that the p.d.f. $g$ satisfies the above backward equation. To show the converse, we technically require regularity conditions to claim the ODE solution is unique, which can often be proved by the contraction mapping theorem. Here, we skip the converse part.
@@
Then, the proof is competed. ∎
B.2 Proof of Theorem 2
@@
This is the HBJ equation in CTMC.
@@ B.4 Proof of Theorem 4
$\underbrace{\frac{dH_{t}(x)}{dt}}_{\text{l.h.s.}}=\underbrace{\sum_{y\neq x}Q^{\theta^{\star}}_{y,x}(t)H_{t}(y)-\sum_{y\neq x}Q^{\theta^{\star}}_{x,y}(t)H_{t}(x)}_{\text{r.h.s.}},\quad p_{ini}=H_{0}(\cdot).$
@@
Finally, we can see that $l.h.s.=r.h.s.$ Furthermore, recalling we have an assumption that $p_{ini}$ is Dirac delta distribution, we clearly have $p_{ini}=H_{0}(\cdot)$. Hence, the statement is proved by the Kolmogorov forward equation.
Appendix C Details of Algorithm
C.1 Straight-Through Gumbel Softmax

### v2

@@ 5.1 Proof Sketch of Theorem 1
$\mathbb{E}_{x_{t:T}\sim P^{\theta^{\star}}}\left[r(x_{T})-\alpha\int_{s=t}^{T}\sum_{y\neq x_{s}}\left\{Q^{\theta^{\star}}_{x_{s},y}(s)-Q^{\theta_{\mathrm{pre}}}_{x_{s},y}(s)+Q^{\theta^{\star}}_{x_{s},y}(s)\log\frac{Q^{\theta^{\star}}_{x_{s},y}(s)}{Q^{\theta_{\mathrm{pre}}}_{x_{s},y}(s)}\right\}ds\mid x_{t}=x\right].$ (9)
This is the expected return starting from state $x$ at time $t$ following the optimal policy. Once the optimal value function is defined, the optimal generator can be expressed in terms of this value function using the Hamilton-Jacobi-Bellman (HJB) equation in CTMC, as shown below.
@@
For $x\neq y$ ($x,y\in\mathcal{X}$), the optimal CTMC generator $Q^{\theta^{\star}}_{x,y}(t)$ can be expressed by the pretrained CTMC generator $Q^{\theta_{\mathrm{pre}}}_{x,y}(t)$ and the value function 9:
$Q^{\theta^{\star}}_{x,y}(t)=Q^{\theta_{\mathrm{pre}}}_{x,y}(t)\exp(\{V_{t}(y)-V_{t}(x)\}/\alpha).$ (10)
@@
The value function 9 follows the Feynman–Kac formulation with respect to the pretrained CTMC distribution $P^{\theta_{\mathrm{pre}}}$ generated by $Q^{\theta_{\mathrm{pre}}}$ as follows
$\exp(V_{t}(x)/\alpha)=\mathbb{E}_{x_{t:T}\sim P^{\theta_{\mathrm{pre}}}}[\exp(r(x_{T})/\alpha)|x_{t}=x]$ (11)
@@
The marginal distribution at time $t$ by the CTMC trajectory in (6) with generator $Q^{\theta^{\star}}(t)$, i.e., $p^{\star}_{t}\in\Delta(\mathcal{X})$, satisfies
$p^{\star}_{t}(\cdot)\propto\exp(V_{t}(\cdot)/\alpha)p^{\mathrm{pre}}_{t}(\cdot),$ (12)
where $p^{\mathrm{pre}}_{t}\in\Delta(\mathcal{X})$ is a marginal distribution induced by pretrained CTMC at $t$.
@@ 5.2 Relation to Classifier Guidance for Discrete Diffusion Models
$\noindent{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\exp(V_{t}(\cdot)/\alpha)=\mathop{\mathrm{argmin}}_{g:\mathcal{X}\to\mathbb{R}}\mathbb{E}_{x_{T}\sim P^{\theta_{\mathrm{pre}}}(x_{T}|x_{t})}[\{\exp(r(x_{T})/\alpha)-g(x_{t})\}^{2}]}}.$ (13)
In practice, while we can’t calculate the exact expectation, we can still replace it with its empirical analog. Alternatively, we can approximate it by using a map from $x_{t}$ to $x_{0}$ in pretrained models following DPS (Chung et al., 2022b) or reconstruction guidance (Ho et al., 2022).
@@
$Q^{\theta^{\star}}_{x,y}(t)=Q^{\theta_{\mathrm{pre}}}_{x,y}(t)\times p_{t}(z|y)/p_{t}(z|x),\quad p_{t}(z|x_{t}):=\mathbb{E}_{x_{t:T}\sim P^{\theta_{\mathrm{pre}}}}[p(z|x_{T})\mid x_{t}].$ (14)
Thus, we can rederive the formula in Nisonoff et al. (2024). Here, we also note that this type of result is commonly referred to as the Doob transform in the literature on stochastic processes (Levin & Peres, 2017, Section 17).
@@ 6 Experiments
Our experiments focus on the design of regulatory DNA sequences for enhancer activity and protein sequences for stability. Our results include comprehensive evaluations, highlighting the ability of DRAKES to produce natural-like sequences while effectively optimizing the desired properties. We also experiment in text generation for minimizing toxicity, detailed in Appendix F.4.
@@ 6.1 Baselines
We compare DRAKES against several baseline methods discussed in Section 2, which we summarize below with further details in Appendix F.1.
@@
  • Classifier-free Guidance (CFG) (Ho & Salimans, 2022). CFG is trained on labeled datasets with the measured attributes we aim to optimize.
@@ 6.2 Regulatory DNA Sequence Design
Dataset and settings. We experiment on a publicly available large-scale enhancer dataset (Gosai et al., 2023), which measures the enhancer activity of $\sim$700k DNA sequences (200-bp length) in human cell lines using massively parallel reporter assays (MPRAs), where the expression driven by each sequence is measured. We pretrain the masked discrete diffusion model (Sahoo et al., 2024) on all the sequences. We then split the dataset and train two reward oracles (one for finetuning and one for evaluation) on each subset, using the Enformer (Avsec et al., 2021) architecture to predict the activity level in the HepG2 cell line. These datasets and reward models are widely used in the literature on computational enhancer design (Lal et al., 2024; Uehara et al., 2024c; Sarkar et al., 2024). Detailed information about the pretrained model and reward oracles is in Appendix F.2.
@@
  • Predicted activity based on the evaluation reward oracle (Pred-Activity). We predict the enhancer activity level in the HepG2 cell line using the reward oracle trained on the evaluation subset. Note that the models are fine-tuned (or guided) with the oracle trained on a different subset of the data, splitting based on chromosome following conventions (but in the same cell lines) (Lal et al., 2024).
@@
  • Approximated log-likelihood of sequences ( App-Log-Lik). We calculate the log-likelihood of the generated sequences with respect to the pretrained model to measure how likely the sequences are to be natural-like. Models that over-optimize the reward oracle generate out-of-distribution sequences and would have a low likelihood to the pretrained model. The likelihood is calculated using the ELBO of the discrete diffusion model in  Sahoo et al. (2024).
Results. DRAKES generates sequences with high predicted activity in the HepG2 cell line, as robustly measured by Pred-Activity and ATAC-Acc (Table 1). The generated sequences closely resemble natural enhancers, as indicated by high 3-mer and JASPAR motif correlations, and a similar likelihood to the pretrained model. These highlight DRAKES’s effectiveness in generating plausible high-activity enhancer sequences. Notably, while DRAKES, without KL regularization achieves higher Pred-Activity, this can be attributed to over-optimization. Despite splitting the data for fine-tuning and evaluation, the sequences remain highly similar due to many analogous regions within each chromosome. However, when evaluated with an independent activity oracle, ATAC-Acc, DRAKES  demonstrates superior performance while maintaining higher correlations and log-likelihood. Ablations on gradient truncation and Gumbel Softmax schedule are in Appendix F.2.
@@
Method Pred-Activity (median) $\uparrow$ ATAC-Acc $\uparrow$ (%) 3-mer Corr $\uparrow$ JASPAR Corr $\uparrow$ App-Log-Lik (median) $\uparrow$ Pretrained 0.17(0.04) 1.5(0.2) -0.061(0.034) 0.249(0.015) -261(0.6) CG 3.30(0.00) 0.0(0.0) -0.065(0.001) 0.212(0.035) -266(0.6) SMC 4.15(0.33) 39.9(8.7) 0.840(0.045) 0.756(0.068) -259(2.5) TDS 4.64(0.21) 45.3(16.4) 0.848(0.008) 0.846(0.044) -257(1.5) CFG 5.04(0.06) 92.1(0.9) 0.746(0.001) 0.864(0.011) -265(0.6) DRAKES w/o KL 6.44(0.04) 82.5(2.8) 0.307(0.001) 0.557(0.015) -281(0.6) DRAKES 5.61(0.07) 92.5(0.6) 0.887(0.002) 0.911(0.002) -264(0.6)
@@ 6.3 Protein Sequence Design: Optimizing Stability in Inverse Folding Model
Dataset and settings. First, we pretrained an inverse folding model based on the diffusion model (Campbell et al., 2024) and the ProteinMPNN (Dauparas et al., 2022) architecture, using the PDB training set from Dauparas et al. (2022). Next, we trained the reward oracles using a different large-scale protein stability dataset, Megascale (Tsuboyama et al., 2023), which includes stability measurements (i.e., the Gibbs free energy change) for $\sim$1.8M sequence variants from 983 natural and designed domains. Following dataset curation and a train-validation-test splitting procedure from Widatalla et al. (2024) (which leads to $\sim$0.5M sequences on 333 domains) and using the ProteinMPNN architecture, we constructed two reward oracles – one for fine-tuning and one for evaluation, that predict stability from the protein sequence and wild-type conformation. Detailed information on the pretrained model and reward oracles is in Appendix F.3.
@@
Results. For inverse protein folding, DRAKES generates high-stability protein sequences capable of folding into the conditioned structure (Table 2). It achieves the highest Pred-ddG among all methods, while maintaining a similar success rate of inverse folding (measured by %(scRMSD$<2$), the percentage of scRMSD smaller than 2) as the pretrained model. Considering both factors, DRAKES significantly outperforms all baseline methods in terms of overall success rate. Note that CFG does not work well for protein sequence design due to limited labeled data, as Megascale includes only a few hundred backbones, making generalization difficult. This is expected, as we mention in Section 2. Further details are provided in Appendix F.1. We provide additional results with pLDDT and sequence diversity, as well as ablation studies on the gradient truncation number and Gumbel Softmax temperature schedule in Appendix F.3.
@@
For validation purposes, we calculate the stability (i.e., Gibbs free energy) of the generated sequences using physics-based simulations (PyRosetta (Chaudhury et al., 2010)) for wild-type protein backbone structures in Figure 2, following (Widatalla et al., 2024). Although all models are conditioned on the same set of protein backbones, different sets of sequences generated by generative methods can lead to significant differences in side chain interactions, which affect folding energies. The results demonstrate that sequences generated by DRAKES are more stable in this validation compared to other baselines. For additional results, refer to Figure 6 in Appendix F.3.
@@ 7 Conclusions
Reproducibility Statement
For the theoretical results presented in the paper, we provide explanations of assumptions and complete proofs in Appendix C. For the proposed algorithm and experimental results, we provide detailed explanations of the algorithm implementations and experimental setup in Section 4.2, Section 6, Appendix E, and Appendix F, and the code and data can be found in https://github.com/ChenyuWang-Monica/DRAKES. For the datasets used in the experiments, we utilize publicly available datasets and elaborate the data processing procedures in Section 6 and Appendix F.
Acknowledgement
CW and TJ acknowledge support from the Machine Learning for Pharmaceutical Discovery and Synthesis (MLPDS) consortium, the DTRA Discovery of Medical Countermeasures Against New and Emerging (DOMANE) threats program, and the NSF Expeditions grant (award 1918839) Understanding the World Through Code.
Appendix A More Related Works
Dirichlet diffusion models for discrete spaces.
Relative trajectory balance for posterior inference. Venkatraman et al. (2024) proposed to use relative trajectory balance to train a diffusion model that sample from a posterior $x\sim p^{\text{post}}(x)\propto p(x)r(x)$. When $r(x)$ is defined as the exponential reward, it can be utilized for reward optimization. However, this approach requires estimation of the normalizing constant term of the unnormalized density. In contrast, we solve a control problem with direct backpropagation. Besides, Venkatraman et al. (2024) models continuous diffusion as a Markovian generative process and is not specifically tailored to discrete diffusion models.
Appendix B Potential Limitations
We have formulated the RL problem, (5), in the context of CTMC. The proposed algorithm in our paper to solve this problem requires reward models to be differentiable. Since differentiable models are necessary when working with experimental offline data, this assumption is not overly restrictive. Moreover, many state-of-the-art sequence models mapping sequences to functions in biology are available today, such as enformer borozi. In cases where creating differentiable models is challenging, we recommend using PPO-based algorithms (Black et al., 2023; Fan et al., 2023) or reward-weighted MLE (Uehara et al., 2024a, Section 6.1).
Appendix C Proof of Theorems
C.1 Preparation
We prepare two well-known theorems in CTMC for the pedagogic purpose. For example, refer to Yin & Zhang (2012) for the more detailed proof. In these theorems, we suppose we have the CTMC:
@@ C.1 Preparation
Here, we prove that the p.d.f. $g$ satisfies the above backward equation. To show the converse, we technically require regularity conditions to claim the ODE solution is unique, which can often be proved by the contraction mapping theorem under certain regularity conditions. Since this is a well-known result, we skip the converse part. If readers are interested in details, we refer to Yin & Zhang (2012).
@@
Then, the proof is completed. ∎
C.2 Proof of Theorem 2
@@
This is the HJB equation in CTMC.
@@ C.4 Proof of Theorem 4
$\underbrace{\frac{dH_{t}(x)}{dt}}_{\text{l.h.s.}}=\underbrace{\sum_{y\neq x}Q^{\theta^{\star}}_{y,x}(t)H_{t}(y)-\sum_{y\neq x}Q^{\theta^{\star}}_{x,y}(t)H_{t}(x)}_{\text{r.h.s.}},\quad p_{\mathrm{ini}}=H_{0}(\cdot).$
Remark 4.
Here, note that $p_{\mathrm{ini}}=\delta(\cdot=\mathrm{Mask})=H_{0}(\cdot)$ holds because we have considered a scenario where the initial is the Dirac delta distribution. When the original initial distribution is stochastic, we will see in Appendix D that we still have $p_{\mathrm{ini}}=H_{0}(\cdot)$.
@@
Finally, we can see that $l.h.s.=r.h.s.$ Furthermore, recalling we have an assumption that $p_{\mathrm{ini}}$ is Dirac delta distribution, we clearly have $p_{\mathrm{ini}}=H_{0}(\cdot)$. Hence, the statement is proved by the Kolmogorov forward equation.
Appendix D Extension with Stochastic Initial Distributions
When initial distributions are stochastic, we need to modify algorithms. Although there are several strategies to take stochastic initial distributions into account (Uehara et al., 2024b; Domingo-Enrich et al., 2024), we consider the strategy of learning the initial distributions (Uehara et al., 2024b).
Hence, we consider the following control problem:
$\theta^{\star}\textstyle$ $=\mathop{\mathrm{argmax}}_{\theta}\underbrace{\mathbb{E}_{x_{0:T}\sim P^{\theta},x_{0}\sim P^{\theta}_{0}}\left[r(x_{T})\right]}_{\textrm{Reward term}}$ (18)
$-\alpha\underbrace{\left\{\mathbb{E}_{x_{0:T}\sim P^{\theta},x_{0}\sim P^{\theta}_{0}}\left[\int_{t=0}^{T}\sum_{y\neq x_{t}}\left\{Q^{\theta_{\mathrm{pre}}}_{x_{t},y}(t)-Q^{\theta}_{x_{t},y}(t)+Q^{\theta}_{x_{t},y}(t)\log\frac{Q^{\theta}_{x_{t},y}(t)}{Q^{\theta_{\mathrm{pre}}}_{x_{t},y}(t)}\right\}dt\right]+\mathrm{KL}(P^{\theta}_{0}\mid p_{\mathrm{lim}})\right\}}_{\textrm{KL term}}.$
Compared to the control problem (5), in our new formulation (18), we parameterize not only generators but also initial distributions. In this case, Theorem 1 still holds.
D.1 Proof
Recalling the definition of value functions:
$V(k,t)=\mathbb{E}_{x_{k:T}\sim P^{\theta}}\left[r(x_{T})-\alpha\int_{t=k}^{T}\sum_{y\neq x_{t}}\left\{Q^{\theta_{\mathrm{pre}}}_{x_{t},y}(t)-Q^{\theta}_{x_{t},y}(t)+Q^{\theta}_{x_{t},y}(t)\log\frac{Q^{\theta}_{x_{t},y}(t)}{Q^{\theta_{\mathrm{pre}}}_{x_{t},y}(t)}\right\}dt\right],$
the optimal initial distribution needs to satisfy
$\mathop{\mathrm{argmax}}_{\theta}\mathbb{E}_{x_{0}\sim P^{\theta}_{0}}[V_{0}(x_{0})]-\alpha\mathrm{KL}(P^{\theta}_{0}\mid p_{\mathrm{lim}}).$
This is equal to the distribution proportional to
$\exp(V_{0}(x_{0})/\alpha)p_{\mathrm{lim}}(x_{0}).$
The rest of the proof holds as in the proof of Section C.4 without any change, referring to Remark 4.
Appendix E Details of Algorithm
E.1 Straight-Through Gumbel Softmax